# Mental Health Impacts of AI Companions

### v1

@@ 6.2.3. Risk Detection and Content Moderation
6.2.4. Privacy, Trust, and Ethical issues
As AICCs become more integrated into users’ emotional and social lives, concerns regarding privacy, trust, and ethical implications intensify. In our qualitative study, participants noted that they shared sensitive personal information with their AICCs, but also expressed concerns about how their data was stored and potentially misused. A few participants mentioned being hesitant to disclose personal details because of worries about privacy violations or data exploitation. As P4 noted, “I was trying to find out immediately after if my conversations are linked to my account… because I didn’t want anything traceable.” We also note the likelihood of an observer effect—users may feel that being observed or monitored could alter their interactions with AICCs (Saha et al., 2024). Together, our findings suggest that risks are not limited to data misuse alone, but also extend to how perceptions of surveillance can shape user disclosure and trust. This highlights the need for robust privacy safeguards, including encryption and transparent data management practices. Users should have full control over their data, with clear options to review, modify, or delete their shared information. Developers must be transparent about data usage, ensuring users understand how their data is stored and whether it may be shared with third parties.
@@ 6.3. Limitations and Future Directions
Our research offers a comprehensive view of psychosocial impacts of AICC use, but has limitations which also inspire interesting future directions. The quantitative component of our study focused exclusively on Replika users, limiting the generalizability of observed psychosocial effects to other AI companion platforms that differ in design, conversational affordances, or social positioning. While the qualitative component included references to Character.ai, Snap’s MyAI, and other platforms, systematic cross-platform comparisons remain an open area for research. Expanding to multiple platforms would enable more robust analyses of how specific design features mediate emotional, behavioral, and cognitive outcomes. Although stratified propensity score matching helped balance observed baseline characteristics between Treatment and Control groups, the study remains observational and cannot establish a true causality. We cannot rule out unmeasured confounds that may have influenced post-treatment outcomes, such as concurrent life events or offline social changes. Experimental or longitudinal designs with finer-grained temporal tracking could strengthen causal claims and clarify the temporal dynamics of psychosocial change. Our triangulation approach, grounded in Knapp’s relational development model (Knapp, 1978), predominantly captured experiences aligning with the “coming together” trajectory (initiating, experimenting, intensifying, bonding). While this lens helped illuminate how benefits and risks evolve over relationship stages, we were unable to observe later-stage dynamics such as relationship stagnation, decline, or disengagement. Future studies should examine these trajectories, as they may reveal different emotional, behavioral, and social consequences of disengagement from AICCs.
@@ 7. Conclusion
In this work, we examined the psychosocial impacts of AI companion chatbots (AICCs) by combining a large-scale quasi-experimental study on social media with semi-structured interviews. We adopted stratified propensity score matching and Difference-in-Differences regression on large-scale Reddit data, to identify both positive effects—such as increased affective expression, readability, and cognitive and social language, as well as negative effects—heightened loneliness and suicidal ideation expressions. Our interviews, analyzed with Knapp’s relationship development model, showed that users often moved through stages of initiation, escalation, and bonding, where AICCs offered emotional validation but also raised concerns about dependency. These findings provide evidence on how AICCs shape wellbeing, highlight the relational dynamics of human–AI companionship, and suggest design directions for supporting healthy boundaries, mindful engagement, and relationship awareness to maximize benefits while reducing risks.

### v2

@@ 6.2.3. Risk Detection and Content Moderation
6.3. Ethical and Privacy Implications
As AICCs become more integrated into users’ emotional and social lives, concerns regarding privacy, trust, and ethical implications intensify. In our interviews, participants described sharing sensitive personal information with their AICCs, but also expressed concerns about how their data was stored and potentially misused. A few participants mentioned being hesitant to disclose personal details because of worries about privacy violations or data exploitation. For example, as P4 noted, “I was trying to find out immediately after if my conversations are linked to my account, because I didn’t want anything traceable.These tensions highlight how perceived opacity and potential surveillance can chill disclosure, even when users experience AICCs as emotionally supportive. From a design standpoint, this underscores the need for robust privacy protections as core safety features, including end-to-end encryption, clearly articulated data-retention policies, and user-facing controls that allow individuals to review, modify, or delete their shared information. Beyond technical safeguards, developers must provide transparent explanations about data practices, ensuring users understand how their data is stored and whether it may be shared with third parties.
Ethical risks also emerge from the growing use of algorithmic inference on linguistic interaction data. In our quantitative analyses, we inferred psychosocial signals such as depression, anxiety, stress, and suicidal ideation from Reddit language. While such measures are validated for population-level research (Yuan et al., 2023a; Saha et al., 2019c), they remain inherently probabilistic: they may misclassify, are context-sensitive, and can reflect training biases. The classifiers and lexicons may not equally capture the language of marginalized groups (Saha et al., 2019b). If AICCs rely on such models to detect distress, some users’ suffering may be under-recognized, while others may be over-flagged. This suggests that AICCs’ design should incorporate participatory and community-based processes to co-define what “risk,” “support,” and “healthy use” look like in human-AI conversations, and to audit inference pipelines for disparate error patterns.
We also emphasize the need for caution in interpreting and applying the findings of this study. While analyses identify psychosocial patterns at scale, they are not diagnostic, should not be applied to individual users, and do not imply that AICC use is inherently beneficial or harmful for any given individual. Such (mis)interpretations can lead to problematic applications—for example, treating elevated loneliness or grief markers as defining characteristics of all AICC users, thereby reinforcing stigmatizing stereotypes about who turns to these systems and why. Such aggregate trends should not be treated as defining characteristics or profiling of AICC users or used to justify stigmatizing narratives, restrictive policy responses, or punitive moderation practices.
Our findings also bear implications for how safety and crisis response are conceptualized in AICC design. Multiple participants described turning to AICCs during periods of heightened distress or loneliness. When an AICC infers self-harm risk from conversation, an ethical tension arises between intervening (e.g., offering crisis resources or escalating to emergency services) and respecting user privacy and autonomy. Overly aggressive or non-consensual escalation may deter candid disclosure, while insufficient response risks reinforcing a false sense of clinical oversight. Addressing this tension requires multi-stakeholder collaboration involving AI developers, clinicians, policymakers, and end-users. From a design perspective, these technologies should clearly communicate their non-clinical status and limits, employ narrowly scoped and transparent crisis protocols, and avoid positioning AI as a substitute for professional care.
@@ 6.4. Limitations and Future Directions
Our research offers a comprehensive view of the psychosocial impacts of AICC use, but also has limitations that inspire interesting future directions. The quantitative component of our study focused exclusively on Replika users, limiting the generalizability of observed psychosocial effects to other AI companion platforms that differ in design, conversational affordances, or social positioning. While the interviews included references to Character.ai, Snap’s MyAI, and other platforms, systematic cross-platform comparisons remain an open area for research. Expanding to multiple platforms will enable more robust analyses of how specific design features mediate emotional, behavioral, and cognitive outcomes.
Our study relies on publicly available social media data and recruitment from online platforms (e.g., Reddit and Facebook), which introduces self-selection and platform sampling biases. Users who post about AICC use on Reddit are likely more engaged, digitally active, and less stigmatized than the broader population of AICC users. As a result, our findings may overrepresent more vocal or highly involved users while underrepresenting individuals who engage more privately, intermittently, or cautiously with these tools. Future work should incorporate multi-platform data sources and complementary recruitment strategies—such as in-app surveys, cross-platform panels, or purposive sampling—to capture a broader range of user experiences. However, such data collection is often only feasible when conducted within proprietary platforms and is not readily accessible to external researchers, underscoring the need for greater transparency and industry-academia collaboration to study broader user populations.
Also, our interview sample reflects demographic skews of English-speaking online communities. Reddit users tend to be younger and more male-leaning (Pew, 2025). Although similar patterns have been reported in public descriptions of Replika’s user base (Layers, 2024), there is no empirical evidence that Reddit participants are representative of the AICC user population. As a result, the psychosocial dynamics observed in our study may not fully generalize to underrepresented groups or to the full diversity of AICC users. Moreover, our recruitment may limit the cultural scope of our findings, as norms around emotional expression, disclosure, stigma, and technology-mediated companionship vary across sociocultural contexts, underscoring the need for future cross-cultural examination of AICC use.
Finally, although stratified propensity score matching helped balance observed baseline characteristics between Treatment and Control groups, the study remains observational and cannot establish a true causality. We cannot rule out unmeasured confounds that may have influenced post-treatment outcomes, such as concurrent life events or offline social changes. Experimental or longitudinal designs with finer-grained temporal tracking could strengthen causal claims and clarify the temporal dynamics of psychosocial change.
@@ 7. Conclusion
In this work, we examined the psychosocial impacts of AI companion chatbots (AICCs) by combining a large-scale quasi-experimental study on social media with semi-structured interviews. We adopted stratified propensity score matching and Difference-in-Differences regression on large-scale Reddit data, to identify both positive effects— such as increased readability, and cognitive and social language, as well as negative effects—heightened depression, loneliness, and suicidal ideation expressions. Our interviews, analyzed with Knapp’s relationship development model, showed that users often moved through stages of initiation, escalation, and bonding, where AICCs offered emotional validation but also raised concerns about dependency. These findings provide evidence on how AICCs shape wellbeing, highlight the relational dynamics of human–AI companionship, and suggest design directions for supporting healthy boundaries, mindful engagement, and relationship awareness to maximize benefits while reducing risks.
Acknowledgements.
We acknowledge the computational resources provided by the Aalto Science-IT project. This research was largely conducted during Yuan’s visiting researcher appointment at the University of Illinois Urbana–Champaign, supported by funding from the Helsinki Institute for Information Technology (HIIT). Renwen Zhang is supported by the Nanyang Technological University Start-up Grant (NAP_SUG 025564-00001).
Appendix
To validate the parallel trends assumption, we visually inspected outcome trajectories for Treatment and Control users during the pre-treatment period. Figure A1 shows log-transformed, smoothed trends across psychosocial, behavioral, and cognitive outcomes. The trajectories of Treatment, Control-VA, and Control-NAI users show no systematic divergence prior to the treatment event, supporting the validity of the DiD design.
(a) LIWC: Affective
(b) Grief: Valence
(c) Grief: Activation
(d) Loneliness
(e) Depression
(f) Anxiety
(g) Stress
(h) Suicidal ideation
(i) Interactivity
(j) Topical Diversity
(k) Posting Frequency
(l) Readability
(m) Complexity
(n) Repeatability
(o) LIWC: Cog. & Percep.
(p) LIWC: Social Context
(q) LIWC: Lexical Density
(r) LIWC: Interper. Focus
(s) LIWC: Temp. Reference
Figure A1. Visualizing parallel trends assumption for DiD regression. The trends are log-transformed with lowess smoothed outcomes trajectories for Treatment and Control groups are shown for one year before and after the treatment or placebo event for three groups: Treatment, Control-LLM, Control-VA, and Control-NAI. Each subfigure shows the outcome trajectory relative to the treatment day (vertical dotted line). The similarity of pre-treatment trends supports the validity of the DiD analysis.
The figure contains multiple line plots showing parallel trends for difference-in-differences regression. Each subfigure presents a specific affective, behavioral or cognitive outcome measured from Reddit posts. Outcomes include affective language, grief valence and activation, loneliness, depression, anxiety, stress, suicidal ideation, interactivity, topical diversity, posting frequency, readability, complexity, repeatability, cognitive and perceptual language, social context, lexical density, interpersonal focus, and temporal references. For each outcome, three groups are compared: treatment, Control-AI, and Control-NAI. Each plot shows log-transformed trajectories for one year before and after treatment, with a vertical dotted line marking the treatment date. Across measures, pre-treatment trajectories are similar between groups, supporting the validity of the parallel trends assumption in DiD regression.